Which critical value is appropriate for a​ 99% confidence level where

nequals=​17,

sigmaσ

is​ unknown, and the population appears to be normally​ distributed?

A.

z Subscript alpha divided by 2zα/2equals=2.583

B.

t Subscript alpha divided by 2tα/2equals=2.898

C.

z Subscript alpha divided by 2zα/2equals=2.567

D.

t Subscript alpha divided by 2tα/2equals=2.921
Answer #1

Solution :

Given that,

sample size = n = 17

Degrees of freedom = df = n - 1 = 17 - 1 = 16

At 99% confidence level the t is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

t /2,df = t0.005,16 = 2.921